OpenStudy (anonymous):

Pleaseeeee help me ASAP f(x)= 1-2(3)^4-2x 1. List the transformations 2. Say where the points (-2, 1/9), (0,1), and (2,9) as well as the horizontal y-asymptote y=o go

OpenStudy (anonymous):

@princemartin I can't see your drawing anymore?

OpenStudy (anonymous):

Is this right? \[f(x) = 1-2(3)^{4-2x}\]

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

1) What does \[f(x) =3^x\] look like? \[f(x) = -1\cdot 3^x\] look like? what does \[f(x) = 3^{-x}\] look like then \[f(x) = 3^{-2x}\] \[f(x)=3^{1-2x}\] then

OpenStudy (anonymous):

1) what does the -2 do in front of the x? 2) what does adding 1 to get 1-2x do? 3) what does multiplying by -2 to do? 4) what does adding 1 do?

OpenStudy (anonymous):

I would answer.. but this hw he hasn;t gone over so i really have no idea nor understand the book. I know it has something to do with shifts and shrinks and all that.

OpenStudy (anonymous):

Ok. I can't give you answers or the mods hammer me. I will do this for a similar problem for you to learn on and then you can do it yourself, is that ok?

OpenStudy (anonymous):

Yeahh thanks!

OpenStudy (anonymous):

Ok if you had \[f(x) = 4^x\] When x was negative you would have something approaching zero and when positive something approaching really big numbers. 

OpenStudy (anonymous):

You know that for this (0,1) is a point since \[4^0 = 1\] and $$4^1=4$$ If you look at $$f(x) = 4^{-x}$$ however you have said all the positive x's are negative and vice versa so you would have: or a mirror over the y axis